# Progressive Binarization with Semi-Structured Pruning for LLMs

### v3

@@ 3.3 Coarse-to-Fine Search
Selecting effective pruning elements is a key challenge in model compression. To address this, we adopt a two-stage strategy. In the coarse stage, we assign pruning ratios to each layer based on its layer importance. In the fine stage, we determine the specific elements to prune within each layer according to the assigned ratios, allowing for more accurate selection of pruning candidates.
Coarse-Stage: Pruning Ratio Allocation.    To estimate the importance of each layer, we adopt a simple criterion that quantifies how much a layer influences the input-output relationship. Layers that induce greater changes between input and output are considered more important, whereas those with limited impact are deemed less important. Specifically, for a given weight layer $\mathcal{W}$, let $\mathcal{X}_{\mathcal{W}}$ and $\mathcal{Y}_{\mathcal{W}}$ denote the input and output, respectively. The importance of the layer is then evaluated as follows:
@@
Inspired by dumitru2024change , we employ cosine similarity to assess the alignment between input and output representations. Based on the computed importance, we rank all layers, with $k_{i}$ denoting the rank of the $i$-th layer. The pruning parameter $N_{i}$ for each layer is determined as follows:
@@
Fine-Stage: Selecting Pruning Elements.    Once the pruning ratio $N/M$ is assigned in the coarse-stage, we proceed to search for the specific pruning elements within each layer. We revisit the impact of pruning on model error. Inspired by OBS hassibi1993optimal , we derive the relationship between pruning and error increase under a second-order approximation, and present the corresponding formulation in Theorem 3.1. The full proof is provided in the supplementary material. Since our objective is to minimize the increase in error caused by pruning, we directly adopt the following score to guide element selection: $s_{q}=\frac{w_{q}^{2}}{\left[\mathbf{H}^{-1}\right]_{qq}^{2}}$. This criterion is theoretically guaranteed to minimize the error increase introduced by pruning under a second-order approximation.Based on the given $N$:$M$ ratio and the corresponding $s_{q}$ values, we select the top $N$ elements with the largest $s_{i}$ values from every $M$ consecutive elements to retain. The other $M-N$ elements are selected for pruning.
@@
After pruning an element $w_{q}$, the pruning-induced error increase $\delta\mathcal{L}$ can be approximated by the following expression: $\delta\mathcal{L}=\frac{1}{2}\frac{w_{q}^{2}}{[\mathbf{H}^{-1}]_{qq}},$ (9) where $[\mathbf{H}^{-1}]_{qq}$ denotes the $q$-th diagonal entry of the inverse Hessian matrix.
@@
4:     $R\leftarrow$ Coarse_Stage_Search$(\mathcal{M},\mathcal{X},\mathcal{TR})$triangleright Allocate layer-wise pruning ratios
@@
6:         $\mathcal{W}_{l}\leftarrow$ Extract_Weights$(\mathcal{M},l)$triangleright Obtain weights of layer $l$
7:         $r_{l}\leftarrow R[l]$triangleright Get pruning ratio for layer $l$
8:         $\mathbf{M}_{\text{p}}\leftarrow$ Fine_Stage_Search$(\mathcal{W}_{l},\mathcal{X},r_{l})$triangleright Generate pruning mask
9:         $\mathcal{W}_{l}^{\prime}\leftarrow$ SPBO$(\mathcal{W}_{l},\mathcal{X},\mathbf{M}_{\text{p}},\tau)$triangleright Stepwise pruning and binarization optimization
10:         $\mathcal{M}^{\prime}\leftarrow$ Update_Layer$(\mathcal{M},l,\mathcal{W}_{l}^{\prime})$triangleright Write updated weights back to model
@@ 3.4 PBS2P Pipeline
Average Bits.    Following BiLLM huang2024billm and STBLLM dong2024stbllm , we similarly divide weights into salient and non-salient groups, consistent with their design. For salient weights, we adopt a residual binarization strategy, allocating two bits to better preserve their expressiveness, as done in both prior works. For non-salient weights, we apply group-wise quantization. In line with STBLLM dong2024stbllm , we further divide the non-salient weights into three groups using two split points, maintaining the same quantization scheme. The weight parameters and additional hardware overhead are as follows:
$\left\{\begin{array}[]{lr}N_{\text{param}}=\left[2\times r_{\text{salient}}+(1-r_{\text{salient}})\right]\times\frac{N}{M},\\ N_{\text{storing}}=2+\dfrac{1}{b_{\text{size}}},\\ \end{array}\right.$ (10)
where $r_{\text{salient}}$ represents the proportion of salient weights, $\frac{N}{M}$ denotes the predefined avrage pruning ratio for the entire model, and $b_{\text{size}}$ indicates the block size used in OBC frantar2022optimal compensation, with 2 bits reserved to mark the division between salient and non-salient weights. Our parameter settings and $\frac{N}{M}$ configuration are identical to our main comparison method STBLLM dong2024stbllm . Specifically, we adopt the same settings of 4:8, 5:8, and 6:8 sparsity as used in STBLLM dong2024stbllm . These correspond to average bit-widths of 0.55, 0.70, and 0.80, respectively.
@@ 4.1 Settings
All experiments are performed using PyTorch paszke2019pytorch and Huggingface paszke1912imperative on a single NVIDIA A800-80GB GPU. Following the work of frantar2022gptq , huang2024billm , and li2024arb , we use 128 samples from the C4 (raffel2020exploring, ) dataset for calibration. Since PBS2P is an efficient PTQ framework, it eliminates fine-tuning, enabling completion through a single process combining binarization and pruning.
Models and Datasets.    We conduct extensive experiments on the LLaMA touvron2023llama1 , LLaMA-2 touvron2023llama2 , and LLaMA-3 dubey2024llama3 families and the OPT family (zhang2022opt, ). To evaluate the effectiveness of PBS2P, we measure the perplexity of LLM’s outputs on WikiText2 (merity2016pointer, ), PTB (marcus1994penn, ), and C4 (raffel2020exploring, ). Moreover, we also evaluate the accuracy for 7 zero-shot QA datasets: ARC-c (clark2018think, ), ARC-e (clark2018think, ), BoolQ (clark2019boolq, ), Hellaswag (zellers2019hellaswag, ), OBQA (mihaylov2018can, ), RTE (chakrabarty2021figurative, ), and Winogrande (sakaguchi2019adversarial, ).
Baselines.    We mainly compare our PBS2P with STBLLM dong2024stbllm , a structural binary PTQ framework designed for compressing LLMs to precisions lower than 1-bit. We compare the results of PBS2P with STBLLM under the same $N$:$M$ settings (e.g., 4:8, 5:8, 6:8). Previous low-bit methods like ARB-LLM li2024arb , BiLLM huang2024billm , GTPQ (frantar2022gptq, ) and vanilla RTN are also selected for comparison.
Table 1: Perplexity comparison of RTN, GPTQ frantar2022gptq , BiLLM huang2024billm , ARB-LLM li2024arb , STBLLM dong2024stbllm , and PBS2P on the LLaMA families. The evaluation results demonstrate the perplexity performance on the Wikitext2 (merity2016pointer, ) dataset across various model sizes.
@@ 4.2 Main Results
We perform a comprehensive comparison of different LLM families (like LLaMA touvron2023llama1 ; dubey2024llama3 ; touvron2023llama2 and OPT zhang2022opt ) with various model sizes. To keep fairness, We follow STBLLM dong2024stbllm to report the average bit-width of all methods, where our methods have the same bit-width as STBLLM.
@@
Furthermore, PBS2P, with a precision of 0.8 bits, outperforms both RTN at 3 bits, GPTQ at 2 bits, BiLLM and ARB-LLM at 1.11 bits, and STBLLM at 0.8 bits in terms of perplexity across all model sizes. Those comparisons show that our PBS2P achieves a better trade-off between bit precision and performance. It is worth noting that PBS2P outperforms GPTQ at 3 bits on LLaMA 1-7B and LLaMA 3-8B. We extend perplexity evaluation to the PTB and C4 datasets. Figure 4 shows the performance of the LLaMA-7B, LLaMA-13B, and LLaMA-2-7B models. PBS2P continues to achieve a leading edge in performance while operating at a relatively lower bit-width compared to other methods.
Results on OPT Family.    We extend our experiments to the OPT family (1.3B to 30B) under sub-1-bit PTQ settings, similar to the setup for LLaMA family. As shown in Table 2, PBS2P continues to outperform STBLLM across most of the models and $N$:$M$ structured binarization configurations. More results are provided in the supplementary material.
@@ 4.3 Zero-Shot Results
Figure 4: BiLLM huang2024billm , ARB-LLM li2024arb , and PBS2P performed on the PTB (marcus1994penn, )and c4 raffel2020exploring datasets, mainly on LLaMA-7B, LLaMA-13B, and LLaMA2-7B, and we found that PBS2P at 0.8 bits outperforms other methods.
@@ 4.4 Ablation Study
Ablation for Calibration Data Size.    Similar to other PTQ-based methods, our PBS2P framework requires only a small calibration set. Table 4f shows the effect of varying calibration size. With the C4 dataset, reducing the sample size to 32 or 64 degrades performance, while increasing to 256 yields minimal additional benefit over 128. These results confirm that PBS2P remains robust with just 128 calibration samples.
Table 4: Ablation study on LLaMA-7B, where all PBS2P is applied an $N$:$M$ sparsity of 4:8. Results are measured by perplexity on the Wikitext2 and C4 datasets. Our results are highlighted in bold.
SPBO Strategy Wikitext2$\downarrow$ C4$\downarrow$
@@
Coarse-Stage Search Metric Wikitext2$\downarrow$ C4$\downarrow$
@@
Prune Type Hardware-friendly Wikitext2$\downarrow$ C4$\downarrow$
@@ 4.5 Time Analysis
As a binary PTQ framework, PBS2P eliminates the need for fine-tuning and retraining. It introduces coarse-to-fine search and stepwise pruning with binarization optimization, which moderately increases the overall computation time. As shown in Table 5, BiLLM huang2024billm completes quantization on LLaMA-7B in 45 minutes, while ARB-LLM li2024arb takes 76 minutes.
@@
5 Discussion
Limitations.  While PBS2P demonstrates strong performance in both accuracy and model compression, it still faces several limitations. First, we have not yet implemented a custom 1-bit kernel that supports the $N$:$M$ sparsity pattern, which limits our ability to accurately evaluate the end-to-end speedup on real hardware platforms. Second, similar to prior binarization methods such as BiLLM huang2024billm , ARB-LLM li2024arb , and STBLLM dong2024stbllm , our approach introduces additional bits to differentiate between salient and non-salient weights. Although this selective binarization strategy helps preserve model performance, it also incurs extra storage overhead, which partially offsets the compression and efficiency gains brought by binarization.
Societal impacts.  Our proposed PBS2P method enables the efficient deployment of large language models on resource-constrained devices, promoting broader access to AI technologies in real-world applications such as mobile assistants, edge computing, and low-power systems. At present, we foresee no negative societal impacts arising from this work.
@@ 6 Conclusion
In this work, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P) for LLM compression. Central to our approach is SPBO, a stepwise semi-structured pruning strategy with binarization optimization, which prunes a subset of elements at each step while jointly optimizing the binarized parameters. SPBO effectively mitigates the combined error from pruning and binarization. Additionally, we introduce a Coarse-to-Fine Search (CFS) strategy to enhance pruning element selection, further improving compression efficiency. Extensive experiments demonstrate that PBS2P outperforms SOTA binary PTQ methods, delivering superior accuracy across various LLM families and evaluation metrics. Our method provides an effective solution for deploying LLMs on resource-constrained devices while maintaining high performance, and offers a novel perspective on combining compression techniques for extreme LLM compression.

### v4

@@ 3.3 Coarse-to-Fine Search
Selecting effective pruning elements is a key challenge in model pruning. To address this, we adopt a two-stage strategy. In the coarse stage, we assign pruning ratios to each layer based on its layer importance. In the fine stage, we determine the specific elements to prune within each layer according to the assigned ratios, allowing for more accurate selection of pruning candidates.
Coarse-Stage: Pruning Ratio Allocation.  To estimate the importance of each layer, inspired by Dumitru et al. (2024), we use cosine similarity to measure the alignment between its input and output representations. We adopt a straightforward criterion that quantifies how strongly a layer shapes the input–output relationship. Layers that induce larger changes are regarded as more important, whereas those with minimal effect are considered less important. Specifically, for a given weight layer $\mathcal{W}$, let $\mathcal{X}_{\mathcal{W}}$ and $\mathcal{Y}_{\mathcal{W}}$ denote its input and output, respectively. The importance of the layer is then computed as follows:
@@
Based on the computed importance, we rank all layers, with $k_{i}$ denoting the rank of the $i$-th layer. The pruning parameter $N_{i}$ for each layer is determined as follows:
@@
Fine-Stage: Selecting Pruning Elements.  Once the pruning ratio $N/M$ is assigned in the coarse-stage, we proceed to search for the specific pruning elements within each layer. We revisit the impact of pruning on model error. Inspired by Hassibi et al. (1993), we derive the relationship between pruning and error increase under a second-order approximation, and present the corresponding formulation in Theorem 3.1. The full proof is provided in the supplementary material. Since our objective is to minimize the increase in error caused by pruning, we directly adopt the following score to guide element selection: $s_{q}=\frac{w_{q}^{2}}{\left[\mathbf{H}^{-1}\right]_{qq}^{2}}$. This criterion is theoretically guaranteed to minimize the error increase introduced by pruning under a second-order approximation. Based on the given NN:$M$ ratio and the corresponding $s_{q}$ values, we select the top $N$ elements with the largest $s_{i}$ values from every $M$ consecutive elements to retain. The other MNM-N elements are selected for pruning.
@@
After pruning an element $w_{q}$, the pruning-induced error increase $\delta\mathcal{L}$ can be approximated by the following expression: $\delta\mathcal{L}=\frac{1}{2}\frac{w_{q}^{2}}{[\mathbf{H}^{-1}]_{qq}},$ (8) where $\mathbf{H}^{-1}]_{qq}$ denotes the $q$-th diagonal entry of the inverse Hessian matrix.
@@
[…]
@@
6:   $\mathcal{W}_{l}\leftarrow$ Extract_Weights(,l)($\mathcal{M}$,l) $\triangleright$ Obtain weights of layer $l$
7:   $r_{l}\leftarrow R[l]$ $\triangleright$ Get pruning ratio for layer $l$
[…]
@@ 3.4 PBS2P Pipeline
Average Bits.  Following BiLLM (Huang et al., 2024) and STBLLM (Dong et al., 2025), we similarly divide weights into salient and non-salient groups, consistent with their design. For salient weights, we adopt a residual binarization strategy, allocating two bits to better preserve their expressiveness, as done in both prior works. For non-salient weights, we apply group-wise quantization. In line with STBLLM (Dong et al., 2025), we further divide the non-salient weights into three groups using two split points, maintaining the same quantization scheme. The weight parameters and additional hardware overhead are as follows:
[…]
where $r_{\text{salient}}$ represents the proportion of salient weights, $\frac{N}{M}$ denotes the predefined average pruning ratio for the entire model, and $b_{\text{size}}$ indicates the block size used in OBC (Frantar & Alistarh, 2022) compensation, with 2 bits reserved to mark the division between salient and non-salient weights. Our parameter settings and $\frac{N}{M}$ configuration are identical to our main comparison method STBLLM (Dong et al., 2025). Specifically, we adopt the same settings of 4:8, 5:8, and 6:8 sparsity as used in STBLLM (Dong et al., 2025). These correspond to average bit-widths of 0.55, 0.70, and 0.80, respectively. This alignment ensures fair, direct comparison under consistent compression budgets.
@@ 4.1 Settings
All experiments are performed using PyTorch (Paszke et al., 2019b) and Huggingface (Paszke et al., 2019a) on a single NVIDIA A800-80GB GPU. Following the work of (Frantar et al., 2023), (Huang et al., 2024), and (Li et al., 2025), we use 128 samples from the C4 (Raffel et al., 2020) dataset for calibration. Since PBS2P is an efficient PTQ framework, it eliminates fine-tuning, enabling completion through a single process combining binarization and pruning.
Models and Datasets.  We conduct extensive experiments on the LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), and LLaMA-3 (Dubey et al., 2024) families and the OPT family (Zhang et al., 2022). To evaluate the effectiveness of PBS2P, we measure the perplexity of LLM outputs on WikiText2 (Merity et al., 2017), PTB (Marcus et al., 1994), and C4 (Raffel et al., 2020). Moreover, we also evaluate the accuracy on seven zero-shot QA datasets: ARC-c (Clark et al., 2018), ARC-e (Clark et al., 2018), BoolQ (Clark et al., 2019), Hellaswag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), RTE (Chakrabarty et al., 2021), and Winogrande (Sakaguchi et al., 2020). This benchmark setup validates our method on both language modeling and reasoning tasks.
Baselines.  We mainly compare our PBS2P with STBLLM (Dong et al., 2025), a structural binary PTQ framework designed for compressing LLMs to precisions lower than 1-bit. We compare the results of PBS2P with STBLLM under the same NN:$M$ sparsity settings (e.g., 4:8, 5:8, 6:8). Previous low-bit methods like ARB-LLM (Li et al., 2025), BiLLM (Huang et al., 2024), GTPQ (Frantar et al., 2023), and vanilla RTN are also selected as baselines for comparison.
Table 1: Perplexity comparison of RTN, GPTQ (Frantar et al., 2023), BiLLM (Huang et al., 2024), ARB-LLM (Li et al., 2025), STBLLM (Dong et al., 2025), and PBS2P on the LLaMA families. The evaluation results demonstrate the perplexity performance on the Wikitext2 (Merity et al., 2017) dataset across various model sizes.
@@ 4.2 Main Results
We perform a comprehensive comparison of different LLM families (like LLaMA (Touvron et al., 2023a; Dubey et al., 2024; Touvron et al., 2023b) and OPT (Zhang et al., 2022)) with various model sizes. To keep fairness, we follow STBLLM (Dong et al., 2025) to report the average bit-width of all methods, where our methods have the same bit-width as STBLLM.
@@
Furthermore, PBS2P, with a precision of 0.8 bits, outperforms both RTN at 3 bits, GPTQ at 2 bits, BiLLM and ARB-LLM at 1.11 bits, and STBLLM at 0.8 bits in terms of perplexity across all model sizes. Those comparisons show that our PBS2P achieves a better trade-off between bit precision and performance. It is worth noting that PBS2P outperforms GPTQ at 3 bits on LLaMA 1-7B and LLaMA 3-8B. We extend perplexity evaluation to the PTB and C4 datasets. Figure 4 shows the performance of the LLaMA-7B, LLaMA-13B, and LLaMA-2-7B models. PBS2P continues to achieve a leading edge in performance while consistently operating at a relatively lower bit-width compared to other methods across diverse evaluation settings.
Results on OPT Family.  We extend our experiments to the OPT family (1.3B to 30B) under sub-1-bit PTQ settings, similar to the setup for LLaMA family. As shown in Table 2, PBS2P continues to outperform STBLLM across most of the models and NN:$M$ structured binarization configurations. More results are provided in the supplementary material.
@@ 4.3 Zero-Shot Results
Figure 4: BiLLM (Huang et al., 2024), ARB-LLM (Li et al., 2025), and PBS2P performed on the PTB (Marcus et al., 1994)and c4 (Raffel et al., 2020) datasets, mainly on LLaMA-7B, LLaMA-13B, and LLaMA2-7B, and we found that PBS2P at 0.8 bits outperforms other methods.
@@ 4.4 Ablation Study
Ablation for Calibration Data Size.  Similar to other PTQ-based methods, our PBS2P framework requires only a small calibration set. Table 4f shows the effect of varying calibration size. With the C4 dataset, reducing the sample size to 32 or 64 degrades performance, while increasing to 256 yields minimal additional benefit over 128. PBS2P remains robust with only 128 calibration samples.
Table 4: Ablation study on LLaMA-7B, where all PBS2P is applied an NN:$M$ sparsity of 4:8. Results are measured by perplexity on the Wikitext2 and C4 datasets. Our results are highlighted in bold.
[…]
@@
[…]
@@
[…]
@@ 4.5 Time Analysis
As a binary PTQ framework, PBS2P eliminates the need for fine-tuning and retraining. It introduces coarse-to-fine search and stepwise pruning with binarization optimization, which moderately increases the overall computation time. As shown in Table 5, BiLLM (Huang et al., 2024) completes quantization on LLaMA-7B in 45 minutes, while ARB-LLM (Li et al., 2025) takes 76 minutes.
@@
Figure 5: Latency comparison (ms) of matrix multiplications across different projection shapes.
4.6 Inference Efficiency
We evaluate inference efficiency using custom matrix multiplication kernels for binarized and semi-structured sparse weights, built on top of the open-source BitBLAS library111 https://github.com/microsoft/BitBLAS. As shown in the figure 5, PBS²P (4:8) consistently achieves the lowest latency across typical LLM projection shapes, with activation sequence length fixed at 2048. Compared to FP16 and other 1-bit baselines like BiLLM and ARB-LLM, PBS²P delivers significant speedups, especially on wider matrices such as 4096×11008 and 11008×4096.
@@ 5 Conclusion
In this work, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P) for LLM compression. Central to our approach is SPBO, a stepwise semi-structured pruning strategy with binarization optimization, which prunes a subset of elements at each step while jointly optimizing the binarized parameters. SPBO effectively mitigates the combined error from pruning and binarization. Additionally, we introduce a Coarse-to-Fine Search (CFS) strategy to enhance pruning element selection, further improving compression efficiency. Extensive experiments demonstrate that PBS2P outperforms SOTA binary PTQ methods, delivering superior accuracy across various LLM families and evaluation metrics. Our method enables efficient LLM deployment on resource-limited devices while preserving performance and unifying compression strategies for extreme model reduction.